OpenStudy (anonymous):

If a, b, c are real numbers such that ac is not equal to zero, show that at least one of the equations 1) ax^2 + bx + c = 0, and 2) -a(x)^2 + bx + c = 0, has real roots.

OpenStudy (anonymous):

You can also show it has non real roots just use values that make them real.

OpenStudy (anonymous):

But, we have to prove it, use discrimant to decide signs, and show the result.

OpenStudy (dumbcow):

discriminant1 = (b^2 -4ac) discriminant2 = (b^2 +4ac) to have real roots one of those must be greater or equal to 0 b^2 -4ac > 0 b^2 +4ac > 0 b^2 > 4ac b^2 > -4ac however this does not prove that either of them are true for any a,b,c

OpenStudy (dumbcow):

wait let me continue b^2 has to be positive, ac can be either positive or negative b^2/ac >4 or b^2/ac > -4 if ac is positive, then b^2/ac > -4 must be true if ac is negative, then b^2/ac >4 must be false

OpenStudy (mani_jha):

This question's already been answered. If ac>0, b^2/ac>0 then b^2>4ac is possible and the first equation may have real roots. In this case, the second equation will have real roots also, since it satisfies b^2/ac>-4 If ac<0, b^2/ac<0 and the first equation wont have real roots. But the second equation will have real roots if -4<(b^2/ac)<0(if b^2/ac lies in between 0 and -4)